# Conditional gene genealogies given the population pedigree for a diploid Moran model with selfing

### v1

@@ 2.1 The pedigree as important partial information of the population
The above population dynamics generates a random graph, which we will call $\mathcal{G}_{N}$, with vertex set $I\times\mathbb{Z}_{+}$ (individuals) and directed edges joining each individual with its parent(s). This random graph $\mathcal{G}_{N}$ is called the pedigree of the population. In the following, we consider the pedigree without specifying outcomes of genetic transmission.
Between consecutive time-steps with two distinct parents, the pedigree has one edge from offspring $(\kappa_{k},k)$ to parent $(\pi_{k,1},k+1)$, one from offspring $(\kappa_{k},k)$ to parent $(\pi_{k,2},k+1)$, and $N-1$ single edges for the same individual from $(j,k)$ to $(j,k+1)$ for $j\in I\setminus\{\kappa_{k}\}$. The portion of the pedigree corresponding to the example above is
@@
Between consecutive time-steps with selfing, the pedigree has a double edges from $(\kappa_{k},k)$ to $(\pi_{k},k+1)$, and again $N-1$ single edges from $(j,k)$ to $(j,k+1)$ for $j\in I\setminus\{\kappa_{k}\}$. The portion of the pedigree corresponding to the example above is
@@
Hence, the offspring always has 2 edges coming out of it which trace upward (towards the past); and all the other $N-1$ individuals have a single edge.
@@
Figure 2: A realization of our diploid Moran process with $N=6$ individuals (left image) and the corresponding pedigree (right image). On the left, two genes $(X_{0},Y_{0})=(2,9)$ are sampled in the present time-step 00, and their lineages are highlighted by the solid dots. Namely, $\{(X_{k},Y_{k})\}_{k=0}^{4}=\{(2,9),(2,9),(2,5),(6,5),(4,4)\}$. These two lineages coalesce in time-step $4$ because $X_{4}=Y_{4}$ but $X_{k}\neq Y_{k}$ for $k=0,1,2,3$.
@@ 2.2 Pairwise coalescence time: unconditional versus conditional on the pedigree
$\tau^{(N,2)}:=\inf\{k\in\mathbb{Z}_{+}:\,X_{k}=Y_{k}\},$ (8)
the (pairwise) coalescence time of the sample of two genes. The left image in Figure 2 shows a realization of the population process in which $\tau^{(6,2)}=4$.
The value of $\tau^{(N,2)}$ is completely determined by the population process. However, knowing the pedigree is not enough to determine the exact value of $\tau^{(N,2)}$, due to Mendelian randomness. For instance, even if we know the pedigree is exactly the one the right in Figure 2 and we know that $X_{0}=2$, we do not know what $X_{3}$ is: $X_{3}$ can be in one of the genes $\{5,6,7,8\}$ with equal probability.
@@ 3 Unconditional distribution of pairwise coalescence time
For $q\in\{\text{diff},\,\text{same},\,\text{coal}\}$, we let $\mathbb{P}_{q}$ be the probability under which the initial state (i.e. the sampling configurations) is $q$. In what follows, we use $\,{\buildrel{d}\over{\longrightarrow}}\,$ and =$\,{\buildrel{d}\over{=}}\,$ to denote convergence in distribution and equal in distribution respectively.
@@
Suppose $\alpha_{N}\rightarrow\alpha\in[0,1]$ as $N\to\infty$. Then $N^{-2}\tau^{(N,2)}$ converges in distribution:
$N^{-2}\tau^{(N,2)}\,{\buildrel{d}\over{\longrightarrow}}\,\begin{cases}0\qquad&\text{under the law }\;\mathbb{P}_{\rm same}(\cdot\,|\mathcal{O}=0)\\ {\rm Exp}\left(\frac{2}{2-\alpha}\right)\qquad&\text{under the laws }\;\mathbb{P}_{\rm diff}(\cdot)\text{ and }\mathbb{P}_{\rm same}(\cdot\,|\mathcal{O}\geq 1)\end{cases}$
where
$\mathcal{O}:=|\{k\geq 1:\,\hat{X}_{k}=\hat{Y}_{k},\,\hat{X}_{k-1}\neq\hat{Y}_{k-1}\}|$
is the total number of time-steps when the two sample lineages transition from belonging to two distinct individuals to belonging to a single individual, and $\hat{X}_{k},\hat{Y}_{k}$ denote the labels of the individuals to which the genes $X_{k}$ and $Y_{k}$ belong.
The convergence to an exponential random variable ${\rm Exp}\left(\frac{2}{2-\alpha}\right)$ in Theorem 3.1, under $\mathbb{P}_{{\rm diff}}$, shows that our Moran model gives the same result as the Wright-Fisher model with partial selfing (Nordborg and Donnelly, 1997; Möhle, 1998) when one averages over the pedigree for a fixed rate of selfing $\alpha\in[0,1]$. Kogan et al. (2023, Lemma 2) recently obtained this result in the broader context of selfing plus recombination under Wright-Fisher reproduction (and averaging over the pedigree).
Two kinds of events are key to our analysis: overlap events where two sample lineages in distinct individuals transition to belonging to the same individual, and splitting events where two sample lineages in the same individual transition to belonging to two distinct individuals. We use $\mathcal{O}$ to denote the number of overlap events. Note that an overlap event results also in coalescence with probability $1/2$. In addition, under $\mathbb{P}_{\rm{same}}$ the two lineages may coalesce before they split, in which case $\mathcal{O}=0$. Under $\mathbb{P}_{{\rm diff}}$ it holds that $\mathcal{O}\geq 1$ almost surely. We prove in Lemma 8.1 that
$\mathbb{P}_{\rm diff}(\mathcal{O}=0)=0\quad\text{and}\quad\mathbb{P}_{\rm same}(\mathcal{O}=0)=\frac{\alpha_{N}}{2-\alpha_{N}}.$ (9)
@@
$\mathbb{P}_{\rm diff}(\tau^{(N,2)}>k)=$ $\,\left(1,0,0\right)\mathbf{\Pi}_{N}^{k}\left(1,1,0\right)^{T}$
$\mathbb{P}_{\rm same}(\tau^{(N,2)}>k)=$ $\,\left(0,1,0\right)\mathbf{\Pi}_{N}^{k}\left(1,1,0\right)^{T}$
Like its Wright-Fisher counterpart (Nordborg and Donnelly, 1997; Möhle, 1998), Theorem 3.1 is a separation-of-timescales result. Direct application of Möhle (1998, Lemma 1, Theorem 1) or its generalization by Möhle and Notohara (2016) is complicated by the extra factor of $N$ in our diploid Moran model, i.e. to go from time in discrete steps to time in units of $N$ generations. In Section 8.3, we instead use our notion of overlap events to prove convergence.
@@ 4 Main results: conditional coalescence times given the pedigree
Our main results are about the conditional distribution of $\tau^{(N,2)}$ given the pedigree and the sampled pair of individuals. We consider a sequence of our diploid Moran model indexed by $N$, under three different regimes for the selfing rate $\alpha_{N}$ in the limit. Our results are summarized in Table 1 and precised in Theorems 4.3 and 4.4. The first corresponds to the model of partial selfing previously studied by others, only excluding cases in which $\alpha_{N}\to 1$ more slowly than $1/N$. The second is what we call limited outcrossing, and happens when $\alpha_{N}\to 1$ at rate $\lambda/N$ for some $\lambda\in(0,\infty)$. The third is when outcrossing is negligible, because $\alpha_{N}\to 1$ faster than $1/N$. As we will show, limited outcrossing interpolates between partial selfing and negligible outcrossing. As $\lambda\to\infty$, it becomes indistinguishable from partial selfing with $\alpha=1$. As $\lambda\to 0$ it becomes indistinguishable from negligible outcrossing.
@@
$\mathcal{A}_{N}:=\sigma\left(\mathcal{G}_{N},\,\hat{X}_{0},\,\hat{Y}_{0}\right).$ (10)
@@
Our main results say that the conditional distributions of $N^{-2}\tau^{(N,2)}$ given $\mathcal{A}_{N}$ converge to different distributions, depending on the sampling configuration (whether the two sampled genes are in the same individual or not) and on the three regimes of the selfing probability. Furthermore, the limiting distributions can retain information of the random pedigree as $N\to\infty$.
Before stating our results rigorously, we first give a quick summary in Table 1 and some illustrations in Figure 3. We are interested in the event ($E_{t}$ in Table 1) that the rescaled conditional coalescence time $N^{-2}\tau^{(N,2)}$ for a sample in state $q\in\{\textrm{diff},\textrm{same}\}$ is greater than a fixed time $t$, namely $\mathbb{P}_{\textrm{q}}(N^{-2}\tau^{(N,2)}>t\mid\mathcal{A}_{N})$ which we denote $\mathbb{P}_{{\rm diff}}(E_{t}\,|\,\mathcal{A}_{N})$ and $\mathbb{P}_{\rm{same}}(E_{t}\,|\,\mathcal{A}_{N})$ in Table 1.
regime selfing rate as $N\to\infty$ $\mathbb{P}_{{\rm diff}}(E_{t}\,|\,\mathcal{A}_{N})$ $\mathbb{P}_{\rm{same}}(E_{t}\,|\,\mathcal{A}_{N})$
partial selfing $N(1-\alpha_{N})\rightarrow\infty$ $e^{-\frac{2}{2-\alpha}t}$ $2^{1-\text{Geom}(1-\alpha)}e^{-\frac{2}{2-\alpha}t}$
limited outcrossing $N(1-\alpha_{N})\rightarrow\lambda\in(0,\infty)$ “random staircase” 0
negligible outcrossing $N(1-\alpha_{N})\rightarrow 0$ $\textbf{1}_{\{{\rm Exp}(2)>t\}}$ 0
Table 1: Summary of results for the conditional coalescence time of a sample of size $n=2$, as $N\to\infty$, where $E_{t}$ denotes the event $\{N^{-2}\tau^{(N,2)}>t\}$ for a fixed time $t$. In the last column, $\text{Geom}(1-\alpha)$ is a geometric random variable $K$ satisfying that $\mathbb{P}(K=k)=\alpha^{k-1}(1-\alpha)$ for $k\in\mathbb{Z}_{+}$. In the middle column, $\textbf{1}_{\{{\rm Exp}(2)>t\}}$ is the indicator of the event $\{{\rm Exp}(2)>t\}$, hence it is a Bernoulli random variable $B$ so that $\mathbb{P}(B=1)=e^{-2t}=1-\mathbb{P}(B=0)$.
From Table 1, we see that under the conditional probability $\mathbb{P}_{{\rm diff}}(\cdot\,|\,\mathcal{A}_{N})$, the limiting survival function of $N^{-2}\tau^{(N,2)}$ changes from deterministic to random, as the selfing rate increases from partial selfing to limited outcrossing. In the partial-selfing regime, the deterministic distribution ${\rm Exp}(\frac{2}{2-\alpha})$ is the same as that for the unconditional case (when we average over pedigrees as in Theorem 3.1). In the limited-outcrossing regime, the limiting survival function is a random random function which we call a “random staircase” for reasons which will become clear. In the negligible-outcrossing regime, the limiting survival function is the random function $\textbf{1}_{\{{\rm Exp}(2)>t\}}$. Note that the expectation of $\textbf{1}_{\{{\rm Exp}(2)>t\}}$, namely $e^{-2t}$, is the same as the unconditional probability $\lim_{N\to\infty}\mathbb{P}_{\text{diff}}(N^{-2}\tau^{(N,2)}>t)$ when $\alpha_{N}\to\alpha=1$. This makes sense because
$\mathbb{E}[\mathbb{P}_{\text{diff}}(N^{-2}\tau^{(N,2)}>t\mid\mathcal{A}_{N})]=\mathbb{P}_{\text{diff}}(N^{-2}\tau^{(N,2)}>t)$ (11)
for all $N\in\mathbb{N}$ and $t\in\mathbb{R}_{+}$. It is always true that the expectation of the conditional limit is the same as the corresponding unconditional limit.
Besides, there is a simple explanation for $\textbf{1}_{\{{\rm Exp}(2)>t\}}$ in the negligible-outcrossing regime in Table 1, by considering the extreme case in which $\alpha_{N}=1$ for all $N$. In this extreme case, there is no outcrossing for the two gene copies of any individual, so there is no splitting event for the lineages of the sampled individuals $\hat{X}_{0},\,\hat{Y}_{0}$. The lineages of the sampled individuals $\{\hat{X}_{0},\,\hat{Y}_{0}\}$ will overlap after a random number of time-steps distributed as a geometric random variable with mean $\frac{N(N-1)}{2}$, hence they overlap after approximately $N^{2}T_{0}$ many generations, where $T_{0}$ is an ${\rm Exp(2)}$ random variable. Upon the overlap, the two sampled genes will stay in the same individual(s) and coalesce after $O(N)$ many time-steps, i.e. coalesce immediately under the $N^{2}$ timescale.
Further, in the case of limited outcrossing, with $N(1-\alpha_{N})\rightarrow\lambda\in(0,\infty)$, we will see that our random walks on the pedigree will converge to a particle system where each ancestral lineage splits at rate $\lambda$ and each pair of ancestral lineages coalesce at rate $2$. In particular, no matter which ancestral lineage our two sample particles belong, they coalesce together at rate $2$, which is what we see in the unconditional case.
Under the conditional probability $\mathbb{P}_{\rm{same}}(\cdot\,|\,\mathcal{A}_{N})$, the limiting survival function of $N^{-2}\tau^{(N,2)}$ changes from random to deterministic (namely 0), as the selfing rate increases from partial selfing to limited outcrossing. Note that in the last column of Table 1, the expectation of the conditional probability
$\mathbb{E}\left[2^{1-\text{Geom}(1-\alpha)}e^{-\frac{2}{2-\alpha}t}\right]=\frac{2-2\alpha}{2-\alpha}e^{-\frac{2}{2-\alpha}t}$
is the same as the unconditional probability
$\lim_{N\to\infty}\mathbb{P}_{\text{same}}(N^{-2}\tau^{(N,2)}>t)=\lim_{N\to\infty}\mathbb{P}_{\text{same}}(\mathcal{O}\geq 1)\mathbb{P}_{\text{same}}(N^{-2}\tau^{(N,2)}>t\mid\mathcal{O}\geq 1)$
according to (9). This makes sense in view of (11).
In Figure 3, we illustrate the limiting behavior of the random, conditional cumulative distribution function (conditional CDF) $t\mapsto\mathbb{P}_{{\rm diff}}(N^{-2}\tau^{(N,2)}\leq t|\mathcal{A}_{N})$ for our three regimes about $\alpha_{N}$.

### v2

@@ 2.1 The pedigree as important partial information of the population
The above population dynamics generates a random graph, which we will call $\mathcal{G}_{N}$, with vertex set $I\times\mathbb{Z}_{+}$ (the $N$ individuals at all time-steps) and directed edges joining each individual with its parent(s). This random graph $\mathcal{G}_{N}$ is the population pedigree. For the moment, let us consider this pedigree without specifying outcomes of genetic transmission.
Between consecutive time-steps with two distinct parents, the pedigree has one edge from offspring $(\varepsilon_{k},k)$ to parent $(\pi_{k,1},k+1)$, one from offspring $(\varepsilon_{k},k)$ to parent $(\pi_{k,2},k+1)$, and $N-1$ single edges for the same individual from $(j,k)$ to $(j,k+1)$ for $j\in I\setminus\{\varepsilon_{k}\}$. The portion of the pedigree corresponding to the example outcrossing event above is
@@
Between consecutive time-steps with selfing, the pedigree has a double edge from $(\varepsilon_{k},k)$ to $(\pi_{k},k+1)$, and again $N-1$ single edges from $(j,k)$ to $(j,k+1)$ for $j\in I\setminus\{\varepsilon_{k}\}$. The portion of the pedigree corresponding to the example selfing event above is
@@
Hence, the offspring always has 2 edges coming out of it which trace upward (towards the past). All the other $N-1$ individuals have a single edge.
@@
Figure 2: A realization of our diploid Moran process with $N=6$ individuals including the genetic transmission events at one locus (left image) and the corresponding pedigree (right image). Also on the left, two gene copies $(X_{0},Y_{0})=(2,9)$ are sampled in the present time-step 00, and their lineages are highlighted by the solid dots. Namely, $\{(X_{k},Y_{k})\}_{k=0}^{4}=\{(2,9),(2,9),(2,5),(6,5),(4,4)\}$. These two lineages coalesce in time-step $4$ because $X_{4}=Y_{4}$ but $X_{k}\neq Y_{k}$ for $k=0,1,2,3$.
@@ 2.2 Pairwise coalescence time: unconditional versus conditional on the pedigree
$\tau^{(N)}:=\inf\{k\in\mathbb{Z}_{+}:\,X_{k}=Y_{k}\},$ (8)
the (pairwise) coalescence time of the sample of two genes. The left image in Figure 2 shows a realization of the population process in which $\tau^{(6)}=4$.
The value of $\tau^{(N)}$ is completely determined by the population process. However, knowing the pedigree is not enough to determine the exact value of $\tau^{(N)}$, due to Mendelian randomness. For instance, even if we know the pedigree is exactly the one on the right in Figure 2 and we know that $X_{0}=2$, we do not know what $X_{3}$ is: $X_{3}$ can be in one of the genes $\{5,6,7,8\}$ with equal probability.
@@ 3 Unconditional distribution of pairwise coalescence time
For $q\in\{\text{diff},\,\text{same},\,\text{coal}\}$, we let $\mathbb{P}_{q}$ be the probability when the initial state (i.e. the sampling configuration) is $q$. We will be interested in the probability the coalescence time exceeds some particular value, or the survival function, which is one minus the cumulative distribution function of the coalescence time. Thus our statements about convergence, in Theorem 3.1 and later Theorems 4.1 and 4.5, are about convergence in distribution.
@@
Suppose $\alpha_{N}\rightarrow\alpha\in[0,1]$ as $N\to\infty$. Then $N^{-2}\tau^{(N)}$ converges in distribution: for any fixed $t\geq 0$
$\mathbb{P}_{q}{\left(N^{-2}\tau^{(N)}>t\right)}\to\begin{cases}\frac{2-2\alpha}{2-\alpha}e^{-\frac{2}{2-\alpha}t}\qquad&\text{when }\;q=\rm same\\[4.5pt] e^{-\frac{2}{2-\alpha}t}\qquad&\text{when }\;q=\rm diff\end{cases}$
The convergence under $\mathbb{P}_{{\rm diff}}$ of the rescaled coalescence time $N^{-2}\tau^{(N)}$ to an exponential random variable ${\rm Exp}\left(\frac{2}{2-\alpha}\right)$ in Theorem 3.1 shows that our Moran model gives the same result as the Wright-Fisher model with partial selfing (Nordborg and Donnelly, 1997; Möhle, 1998) when one averages over the pedigree for a fixed probability of selfing $\alpha\in[0,1]$. Kogan et al. (2023, Lemma 2) recently obtained this result in the broader context of selfing plus two-locus recombination under Wright-Fisher reproduction (and averaging over the pedigree).
The proof of Theorem 3.1 is given in the Appendix, Section 7.1. Two kinds of events are key to our analysis: overlap events where two sample lineages in distinct individuals transition in one time-step to belonging to the same individual, and splitting events where two sample lineages in the same individual transition to belonging to two distinct individuals. We note that these correspond, respectively, to the ‘parent-sharing’ events and ‘separation’ events in the two-locus unconditional model of Uyenoyama (2024). Splitting events occur at outcrossing events, whereas overlap events may be due either to selfing or outcrossing.
@@
$\mathbb{P}_{\rm diff}(\tau^{(N)}>k)=$ $\,\left(1,0,0\right)\mathbf{\Pi}_{N}^{k}\left(1,1,0\right)^{T}\text{ and }$
$\mathbb{P}_{\rm same}(\tau^{(N)}>k)=$ $\,\left(0,1,0\right)\mathbf{\Pi}_{N}^{k}\left(1,1,0\right)^{T}$
exactly in discrete time. Like its Wright-Fisher counterpart (Nordborg and Donnelly, 1997; Möhle, 1998), Theorem 3.1 is a separation-of-timescales result. Direct application of Möhle (1998, Lemma 1) or its generalization by Möhle and Notohara (2016) is complicated by the extra factor of $N$ in our diploid Moran model, i.e. to go from time in discrete steps to time in units of $N$ generations. Kroumi and Lessard (2015, Proposition 1) have extended Möhle’s lemma to the Moran model. In Section 7.1, we instead use our notions of overlap and splitting events to prove convergence.
@@ 4 Main results: conditional coalescence times given the pedigree
Our main results are about the conditional distribution of $\tau^{(N)}$ given the pedigree and the sampled pair of individuals. We consider a sequence of our diploid Moran model indexed by $N$, under three different regimes for the selfing probability $\alpha_{N}$ in the limit. The first corresponds to the model of partial selfing previously studied by others, but holds only when $\alpha_{N}\to 1$ more slowly than $1/N$. The second is what we call limited outcrossing, and happens when $\alpha_{N}\to 1$ at rate $\lambda/N$ for some $\lambda\in(0,\infty)$. The third is when outcrossing is negligible, because $\alpha_{N}\to 1$ faster than $1/N$. As we will show, limited outcrossing interpolates between partial selfing and negligible outcrossing. As $\lambda\to\infty$, it becomes indistinguishable from partial selfing with $\alpha=1$. As $\lambda\to 0$ it becomes indistinguishable from negligible outcrossing.
@@
$\mathcal{A}_{N}:=\sigma{\left(\mathcal{G}_{N},\,\hat{X}_{0},\,\hat{Y}_{0}\right)}$ (9)
@@
which by definition contains all the information relevant to the distribution of the coalescence time. Our main results (Theorems 4.1 and 4.5) say that the conditional distributions of $N^{-2}\tau^{(N)}$ given $\mathcal{A}_{N}$ converge to different distributions, depending on the sampling configuration (diff or same) and on the three regimes of the selfing probability. Furthermore, the limiting distributions can retain information of the random pedigree even as $N\to\infty$.
We now state our main results for coalescence times in Theorems 4.1 and 4.5.
Theorem 4.1.
For any $t\in\mathbb{R}_{+}=[0,\infty)$, as $N\to\infty$,
$\mathbb{P}_{{\rm diff}}{\left(N^{-2}\tau^{(N)}>t\,|\,\mathcal{A}_{N}\right)}\to\begin{dcases}e^{-\frac{2}{2-\alpha}t}&\quad\text{if}\quad N(1-\alpha_{N})\rightarrow\infty\;\text{ and }\,\alpha_{N}\rightarrow\alpha\in[0,1]\\[2.25pt] \mathbb{P}(T_{\lambda}>t\,|\,G_{\lambda})&\quad\text{if}\quad N(1-\alpha_{N})\rightarrow\lambda\in\mathbb{R}_{+}\\[2.25pt] \mathbbm{1}_{\{{\rm Exp}(2)>t\}}&\quad\text{if}\quad N(1-\alpha_{N})\rightarrow 0\end{dcases},$
where $G_{\lambda}$ is the ancestral graph, defined in Definition 4.3, and $T_{\lambda}$ is the coalescence time for a pair of independent simple random walks on $G_{\lambda}$, defined in Definition 4.4.
The proof of Theorem 4.1 is given in the Appendix, Section 7.2.
Remark 4.2.
This result contrasts with the asymptotic behavior of the marginal or unconditional probability $\mathbb{P}_{{\rm diff}}{\left(N^{-2}\tau^{(N)}>t\right)}$ in Theorem 3.1. For example, assuming that $\alpha_{N}=1$ and assuming that $\alpha_{N}\to 1$ at rate $N^{-2}$ lead to the same result in Theorem 3.1 but correspond to negligible outcrossing and partial selfing, respectively, in Theorem 4.1.
Theorem 4.1 offers a description of the conditional distribution of the coalescence time $\tau^{(N)}$ for a sample of two gene copies from different individuals in a population of size $N$ given the pedigree. Simulations of the CDFs for five pedigrees under each of the three cases in Theorem 4.1 are shown in Figure 3. Theorem 4.1 says that the law of $N^{-2}\tau^{(N)}$ for two lineages in different individuals is robust to the structure of the pedigree so long as the rate at which two lineages in the same individual undergo a splitting event, $(1-\alpha_{N})N^{-1}$, is much greater than the rate at which two sample lineages in distinct individuals coalesce, $N^{-2}+O(N^{-3})$, specifically that the ratio of the former over the latter, approximately $N(1-\alpha_{N})$, tends to infinity. This robustness for the case of partial selfing is simply that the conditional limit agrees exactly with the marginal limit in Theorem 3.1.
If, on the other hand, the rate of coalescence is comparable to or dominates the rate of splitting of an individual, as in the cases of limited outcrossing and negligible outcrossing, Theorem 4.1 demonstrates the non-robustness of the “diff” pairwise coalescence time to the pedigree. That is, the survival function of the conditional pairwise coalescence time does not converge to its mean. Indeed, in this case the rates at which lineages split backwards in time is then comparable to the rate at which they coalesce. The subgraph of the pedigree which encodes the possible trajectories of the sample lineages takes the shape of an ancestral graph $G_{\lambda}$ as the population size $N$ goes to infinity, and the ancestries of the two sample lineages correspond to coalescing random walks thereon with coalescence time $T_{\lambda}$. Correspondingly, $N^{-2}\tau^{(N)}$ conditional on the pedigree asymptotically takes the distribution of $T_{\lambda}$ conditional on the ancestral graph.